Fermi National Laboratory
 

Accelerator Update for September 22, 2006 - September 25, 2006

 

Friday September 22

The day shift began with Operations monitoring store 4967 and a Recycler stash and the stack, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beam, with the Antiproton Source (Pbar) having stacking problems, with the Linac, Booster, Main Injector (MI), and Recycler operating normally, and with NuMI and MiniBooNE taking beam.

Stacking resumed at 8:25 AM.

At 9:15 AM, Operations reported that store 4967 aborted; there was no quench. A CAMAC link went down. The controls experts investigated and made repairs.

Operations put the Tevatron (TeV) through a dry squeeze at 9:39 AM.

Operations put the TeV through a wet squeeze at 1 PM.

Operations put the TeV into shot setup at 1:22 PM.

Operations began preparing the H- Source Dome at 2:43 PM for an access.

The H- Source access ended and stacking resumed at 3:02 PM.

Operations established store 4970 at 3:31 PM with an initial luminosity of 167.9E30.

Saturday September 23

The midnight shift began with Operations monitoring store 4970 (current average luminosity of 65.69E30) and the stash (97.8E10), with Pbar stacking, and with NuMI and MiniBooNE taking beam.

At 10 AM, Operations reported that the Pbar target had stopped rotating. An operator contacted experts. Stacking resumed at 10:13 AM, but at a reduced intensity.

Operations halted stacking at 1:32 PM in preparation for an access to work on the Pbar target.

At 4:04 PM, a Linac RF station (KRF6) went down and would not reset. A Linac expert worked on the station. NuMI and MiniBooNE resumed taking beam at 4:17 PM.

At 4:37 PM, a Linac expert reported that KRF6 was having regulation problems.

Operations terminated store 4970 at 5:03 PM. TeV RF experts began work on a RF station (TRF1).

Pbar experts completed their target work. Stacking resumed at 5:24 PM.

Operations put the TeV into shot setup at 6:58 PM.

At 8:20 PM, Operations reported that KRF6 had tripped off four times in the last 20 minutes. Operators contacted an expert.

Operations established store 4972 at 9:26 PM with an initial average luminosity of 162.3E30.

At 10:38 PM, Operations reported that KRF6 was causing more problems. Linac experts worked on the station.

Sunday September 24

The midnight shift began with Operations monitoring store 4972 and the stash, but Linac was off due to KRF6 problems.

Linac expert reported that KRF6 was back in service, but serious problems remained. Pbar resumed stacking and NuMI and MiniBooNE resumed taking beam at 7:03 AM.

At 7:27 AM, KRF6 tripped off. Experts investigated.

TeV experts began Tevatron Electron Lens studies (TEL2) at 9:16 PM.

At 10:40 PM, the Pbar target stopped rotating again. Operations reduced the intensity and resumed stacking at 10:54 PM.

Monday September 25

The midnight shift began with Operations monitoring store 4972 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam.

Operation prepared the Pbar vault at 6:11 AM for an access to work on the target.

Operations terminated store 4972 at 6:35 AM.

D0 accessed their collision hall around 6:53 AM.

Operations prepared the TeV F sector at 7 AM for an access to investigate a kicker magnet.

The Plans for Monday

The plans for today are to establish a store as soon as possible and finish the Pbar target work.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 09/25/2006    email Fermilab